OpenStudy (anonymous):

Which expression is equivalent to 3xy + 10xy - 5x + 2x - 2yx? A) 15xy + 7x B) 13xy - 7x - 2xy C) -8xy D) -3x + 11xy

pooja195 (pooja195):

Combine like terms 3xy + 10xy - 5x + 2x - 2yx<----yx is the same thing as xy (3xy+10xy+−2xy)+(−5x+2x) <---do the math to get the answer :)

OpenStudy (anonymous):

i keep getting it wrong @pooja195

pooja195 (pooja195):

What did you get?

OpenStudy (anonymous):

-27xy

pooja195 (pooja195):

(3xy+10xy+−2xy)+(−5x+2x) 3+10 = ?

OpenStudy (anonymous):

13

pooja195 (pooja195):

(3+10−2)+(−5+2) <---solve this 13-2= ?

OpenStudy (anonymous):

11

pooja195 (pooja195):

good so we have 11xy -5+2= ??

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

so its D?

pooja195 (pooja195):

Yep

OpenStudy (anonymous):

thx

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)